OpenStudy (anonymous):

Simplify the expression, 4xy^2 0^4(-y)^-3 Just this one question!!!

OpenStudy (anonymous):

Help? I know the answer is 0, I just need help on how it got there, please help!

OpenStudy (anonymous):

\[\huge 4x \times y^2 \times 0^4 \times (-y)^{-3}\] \[\huge 4x \times y^{2-3}\times 0 = \huge 4x \times y^{-1}\times 0 =0\]

OpenStudy (anonymous):

I can't make it straight but I sort of got what you typed! Thanks! I see!

OpenStudy (anonymous):

So zero can cancel out both numbers and variables?

OpenStudy (anonymous):

@dpasingh ?

OpenStudy (anonymous):

@Duizhang✨ Any nuber or variable multiplied by zero is zero, so is inthe case of your question.

OpenStudy (anonymous):

Thank you! ^^